40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 4 18:59:39 2018, gautam, Update, SUS, Stack measurement ongoing 
    Reply  Mon May 7 15:27:28 2018, gautam, Update, SUS, Stack measurement expectation 40m_TMstack.JPG
Message ID: 13815     Entry time: Fri May 4 18:59:39 2018     Reply to this: 13821
Author: gautam 
Type: Update 
Category: SUS 
Subject: Stack measurement ongoing 

[SV,KA,RXA,GV]

The stack weight measurement is going on at EX. ETMX watchdog is shutdown. Area is off limits over the weekend until the test is finished.


Not related to this work, but the dog clamps used on the EX table have to be re-positioned such that the clamping force is better distributed. The 2" beam splitter mount used to pick off a portion of the EX NPRO beam to the fiber has to be rotated. Also, there was a M6.9 EQ in Hawaii while we were doing this work it seems..

ELOG V3.1.3-